FEL Upgrade Project Weekly Brief – August 12-16, 2002




Work on the crucial ball cathode support tube was completed this week. See the link below for an image of the prepared components. The high voltage power supply for the gun also made significant assembly progress.


H. F. Dylla spent 2.5 days at a Director’s Council offsite working out-year program planning and priorities for the laboratory.

Planning continued for presentations next week to Ray Orbach, DOE Director of the Office of Science.

We thank the ONR for their continuing efforts to keep our accounts in the black by completing delivery (plus some) of the remaining FY02 funds due. We are preparing July earned value reports at this time.

WBS 4 (Injector):

The work with W&M on the ion implantation of the ball cathode support tube concluded successfully this week. A close examination revealed that the tube is indeed free of arcing marks. A total of six runs were performed to cover the entire surface of the tube with SiO2.

A SS sample will be implanted and field emission scanned for testing up to 60 MV/m (the highest field region on the tube will be about 12 MV/m). Pictures of the implanted support tube and the ball cathode can be seen at http://laser.jlab.org/gun/. The fixturing for welding the support tube to its flange was fabricated.

Gun HVPS - The 3 stacks and center toroids for the HVPS were installed this week. The template to drill the large bottom toroid was made. It will be drilled and installed next week. The frame to hold the HVPS electronics on the tank was installed. The castor wheel on the tank door was changed and the castor re-shimmed to bring the door's bolt holes into alignment. The missing SF6 piping will be installed next week.

WBS 5 (SRF):

Plans to measure HOM damping in a 7 cell cavity continue. In order to get realistic measurements of the HOM Qs we intend to test a seven cell cavity at temperature in a cryogenic dewar with both the high power and HOM couplers attached. This test can be initiated within two weeks or less after the required dewar fit up hardware is machined. Copper cavity tests are also continuing to determine required placement of the HOM loads for optimal coupling.

WBS 6 (RF):

Quarter HVPS - After Unit 3 was hi-potted to 40 kV, the 480 VAC was connected. The unit operated to 27 kV under no-load conditions which is the maximum voltage available from the 70% tap position. We have experienced faulty klystron body current, arc, and crowbar trips that occur when the heater relay makes an adjustment. It appears to be caused by the +24 V power supply being overloaded. A power supply with more current capacity is being tried.

Quarter Klystrons - Hoses to provide LCW to the Circulator, RF Load, and HOM filter for unit 3 were made this week. The other 2 klystron carts will be worked also next week. Shorts will be installed today.

Zone 4 - The RF waveguide connections to the cryomodule are being made by SRF today. The balance of the waveguides will be connected when the parts are available. Flexible couplings are on order and the half-height sections are being made by the Machine Shop.


WBS 8 (Instrumentation):

DLPC - The new EPICS screen for rep-rate and pulse width control have gone through the first wave of testing. It is still in the development stage and is currently being revised. Screen-shot at: http://laser/systems/dlpc/upgrade2002/dlpc-newscreen-rev1.jpg.

BLMs - The manufactured F0151 (front) card is in-house and an assembly package for the set of 6 boards is being prepared for component placement next week. Issues surrounding the software upgrade to the high voltage power supply (LeCroy) and related screens have been addressed the new screens are under development.

MPS - Trent Allison from EEI&C is helping out on the design of the Machine Mode/Beam Mode generator cards. In discussions this week concerning the layout and packaging of the system were are able to proceed with a design that does not need two separate boards for these functions. A 25-pin "D" ribbon cable will straddle the input boards and the gate arrays for each board will be mapped to all of the pins (20 not the 5 ground pins) such that each board will have three pins to 'talk' on and will 'listen' to the rest. The boards will be in-circuit programmable via a 10 pin header on the front panel.

The design review was held for the large magnet power supply installation and switch configuration there were very minor comments - good job done by the EEDC group, thanks! The old CAMAC crate and electronics for the magnet interface have been removed from Zone 5 to facilitate the effort on the magnet communication interface upgrade. This leaves only the RF system in CAMAC. The RS-485 communication cables have been pulled into place for the trim magnet utility chassis. System and installation drawings are being updated as work progresses for this system. Cables were pulled in Zone 5 for the Magnet Thermal Interlock chassis. The 4000 feet of 2 ga. cable is here for inject, extract, and octa-sextapole magnets, the student will enjoy pulling it in over the weekend.

Further progress has been made with the PC/104 board. The students were able to get EPICS to compile correctly under linux on the desktop, and compile an example application (i.e. was able to get an epics IOC shell prompt on the PC). The example code (or any application that we develop) can be copied directly off the PC to the PC/104 board and will run without modification. The PC/104 will serve as a low (~no) cost alternative to the VME based data acquisition solution. The first application is to monitor the optical cavity mirror thermocouples.

Great progress has also been made with the automated emittance measurement project. A new video capture card was selected that will allow maximum usage of the Dell PC for video processing. The one we were using is already obsolete - shelf life worse that fruit. The new card is an Osprey-100. It has three composite video inputs and one s-video input on a single PCI card. The spot size analysis code is being reworked to use multidimensional array, rather than a single dimension array. This will aid in data manipulation and analysis. The image captured is now moved into a multidimensional array, and can be saved. The summation and deletion (background subtraction) functions will be patched up to work the same for consistency. Some time was spent figuring out how to properly pass a multidimensional array with C. This design could also be used for control of the optical cavity mirror heaters. These mirror heaters are to be controlled base on a HeNe spot size.

WBS 9 (Transport):


Optical Chicane Dipoles (GW)

• The remaining 6 cores of the second batch of 6 Magnets for UV passed gap checking and will continue to further processing.

Injector Dipoles (DU/DV)

• The measurements of the first GV verified that the scheme of biasing the core field with a trim coil allows the central field to be made flat to within ± one part in 10,000, a very difficult feat for this "all ends" magnet.

- The field flatness comes with a price however. The main current is 9 amperes higher than the field in the GU partner magnet in the string. The GU will require a shunt that had not been anticipated.

- In addition, both the field flatness and central field are very dependent on trim coil ramp rate. For the QX Quadrupoles, we achieved uniformity with a bang-bang (as fast as you can) current setting method. However, for these trim circuits, the sensitivity of the main fields and field flatness are much greater at rapid turn. We will not be sure that rapid turn on the test stand and rapid turn on in the FEL enclosure will be the same. Instead we will use a defined slow ramp rate on the test stand. For the FEL enclosure we are working on a software fix for these specific trim cards to ramp them at the identical rate. This should be possible with the newly installed, control that has eliminated scanner cards.

• The gluing of shim materials to GU cores passed after examining the detailed records o of the individual measurements and GU magnets were assembled.

• The remaining 5 GVs were not shipped pending resolution of supposed claims for work outside of the scope contract. The resolution was not helped by the fact the Wang NMR has not used any of our forms and procedures for resolving claims but has rather loaded all supposed problems on three contracts into one document with no reference to our documentation. We are working the issue.

Arc 180 Degree Dipoles (GY)

• GY Coil Potting fixture remains in fabrication.

Arc Bend, Reverse Bend Dipoles (GQ, GX)

• Process Equipment Co. (PECo) shipped the four GQ cores to Wang NMR and they arrived in Livermore today. Because the building entrance can't accept the 44 ft. trailer to off-load the magnets, a local machinery mover (familiar with the problem) was employed as a transfer agent.

• At Wang NMR, potting fixtures remain in fabrication for the GQ and GX coils.


Trim Quad (QT)

• The movable traverse mechanism the quad test stand was improved with a hand crank lead screw during the week to improve the quality of the Trim Quad testing.

Sextupole (SF)

• DULY Research continues finishing the Sextupole drawings per our mark-ups.

• We have decided to not use the 6.6 inch thick plate we originally bought for material for the Sextupole project. Rather, we will use the much thinner spare plates that were bought for the dipole cores to obtain the 4 inch thick pieces. This plan eliminates a lot of "hogging" machining. It also uses up a batch of disparate pieces (but still from one heat) that would be difficult to utilize in any other way.

Octupole (OT)

• DULY stopped design of the Octupole while they finish the Sextupole.

Beam Line and Vacuum

• For the Arc Chamber job, Master Machine reports delays in their cutting the major parts that could add a month to the schedule. We are working the issue.

• The JLab Shop completed the full penetration welding of the X Chamber. We are examining the welds with a bore scope as we did for the IR Demo.

• The JLab Shop continues fabrication of the four chambers for the optical chicane region.

• Design of the regions around the wiggler is 90 % complete.

• The extraction chamber design is nearing the checking stage.

• The stands for sextupoles and trim quads are in fabrication.

• Girder assembly continued

• Low Conductivity Water piping installation continued in the FEL enclosure while additional stand were installed and the alignment crew continued stand placement.

WBS 11 (Optics):

This week's mechanical work on the optical cavity primarily was spent on the OCMMS feedback implementation, and building the support frame for the HEPA filtered enclosure the optical cavity assemblies will be built under when they return from cleaning. The Machine Shop also refinished one of the brackets that connects the leadscrew nut to the rest of the cavity hardware. It was checked on the CMM and is now in spec. Next week we will reassemble and test the hardware with the remanufactured part. We held design meetings to review the mounting brackets for the optical cavity water cooling tubes, and on the optical transport turning cassettes. We also met on the design of the collimator. We received notification that the award for the remaining linear translators and the high vacuum LVDTs were awarded.

We are trying some other remediation techniques to improve the power stability of our frequency-locked ultrafast laser. We'll begin taking more phase noise measurements late this afternoon (8/16).

Other activities:

We assisted in verifying the performance of a new injector drive laser (for CEBAF), since we are considering purchase of the advanced drive laser from the same vendor. It performed quite nicely. We also spent some time with potential vendors of the laser we'll use with the FEL external cavity (the pulse stacker).